EE 308 Microcontrollers Syllabus: Spring 1996

Instructor:  Bill Rison
Office:  Workman 39
Office Hours: M W 11-12, Tu 1-2, and by appointment

E-Mail:  rison@ee.nmt.edu
URL:  http://www.ee.nmt.edu/~rison

Phone:  835-5486
Texts:  Microcontroller Engineering

Gene H. Miller

Motorola Data Manuals

General Outline (Tentative):

• Introduction to the 68HC11 Microcontroller

• Binary and Hexadecimal Numbers

• Assembly Language Programming

• C Language Programming

• Introduction to 68HC11 Internal Peripherals – The 68HC11 Timer Subsystem

– Interrupts using the Timer Subsystem • The 68HC11 Expanded Mode

– Address and Data Buses and Timing

– Adding Memory and External Peripherals – Using the 6821 PIA for Parallel I/O

• More 68HC11 Internal Peripherals – The A/D Converter Subsystem

– The Serial Communication Interface

Homework: 15%
Quizes:  10%
Exams:  50%
Final:  25%

Late homework will be accepted, but the score will be decreased by 10% for each class period it is late. Homework more than one week late will not be accepted. There will be four exams.